OpenStudy (anonymous):

Find the number of ways to listen to 4 CDs from a selection of 8 CDs. 1680 336 70 50

OpenStudy (anonymous):

does the order matter?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

then 8P4=8*7*6*5=1680

OpenStudy (anonymous):

u know the reason?

OpenStudy (anonymous):

we can choose the first cd in 8 ways, the next in 7 ways, the next in 6 ways and the last in 5 ways.. so, total 8*7*6*5=1680